# Gravity as a topological gauge theory

### v3

@@ 2.3 Associated action
The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the $\mathfrak{g}$-valued curvature $\bar{\Omega}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+2k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}$.
By combining linearly the Euler number ($\mathdutchcal{E}$ (5)) with the two Pontrjagin numbers ($\mathdutchcal{P}$ (7)) associated to $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}$, one can build the action (for $e,r,y$ real numbers):
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(\bar{\Omega}_{\mathfrak{h}})=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})+yP(\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}\left(\tfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi}\right)+rP(\bar{\Omega})+yP(\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
@@
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{8k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;b}d^{4}x+\dfrac{96k^{2}}{\ell^{4}}\sqrt{|\tilde{g}|}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)(R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;d}\varepsilon^{cd}_{~{}~{}ab}d^{4}x)-\dfrac{2rk}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}$ (1)
@@
As mentionned in [7, p.192] there exists an isomorphism between elements of the Lorentz Lie algebra $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$ that can be written as $\Phi_{\varepsilon}:\mathfrak{so}(n-1,1)\rightarrow\mathfrak{spin}(n-1,1)$. Its action on the Lie algebra generators of $\mathfrak{so}(n-1,1)$ is:
$\Phi_{\varepsilon}(J_{rs})=\dfrac{1}{4}\bar{e}_{a}J_{rs}e_{b}\gamma^{a}\gamma^{b}=\dfrac{1}{4}(J_{rs})_{ab}\gamma^{a}\gamma^{b}=\dfrac{1}{2}\eta_{ar}\eta_{bs}\gamma^{a}\gamma^{b}.$ (2)
From the covariant derivative on the space of spinors $D_{\mathfrak{h}}=d+\Phi_{\varepsilon}(\varpi_{\mathfrak{h}})=(\partial_{\mu}+\dfrac{1}{4}\varpi_{\mathfrak{h}\>\,ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$ associated to $\varpi_{\mathfrak{h}}$ and $h^{\varepsilon}$ one can construct the following matter action: [^7] Since the metrics at the level of the Lie algebra and on $\mathscr{M}$ are always defined up to the factors $\kappa$ and $\zeta$, we set $\kappa^{2}\zeta=\tfrac{\ell^{3}}{216k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [7, p.197].
$S_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=\mathrm{Re}S_{D}$ (3)
where $\mathrm{Re}S_{D}$ corresponds to the real part of the Dirac action as described in [7, p.197].
We are now in position to identify the different terms of $S_{G}[A,\beta]$ above with the action presented in [8] which, in our notation, reads:
$S^{\prime}[\varpi]$ $=\int_{\mathscr{M}}\Biggl{(}\biggl{(}\overbrace{\dfrac{\alpha_{1}\ell^{2}}{2}R^{ab}\wedge\star\xi_{ab}+\dfrac{\alpha_{2}\ell^{2}}{2}R^{ab}\wedge\xi_{ab}}^{Holst}+\overbrace{\alpha_{3}R^{ab}\wedge R_{ab}}^{Pontrjagin}$ (4)
$+\overbrace{\alpha_{4}R^{ab}\wedge\star R_{ab}}^{Euler}+\overbrace{\alpha_{5}\bigl{(}T^{a}\wedge T_{a}-\dfrac{\ell^{2}}{2}R^{ab}\wedge\xi_{ab}\bigr{)}}^{Nieh-Yan}+\overbrace{\dfrac{\alpha_{6}\ell^{4}}{4}\varepsilon_{abcd}\xi^{ab}\wedge\xi^{cd}}^{Cosmological~{}constant}\biggr{)}\Biggr{)}$
@@
In this case from the first Bianchi identity $D\bar{\Omega}_{\mathfrak{h}}=0$, and by neglecting boundary terms, the equations of motion for $A$ and $\beta$ require respectively a null spin density and a null energy-momentum tensor and thus turns out to be physically irrelevant.
While for general mutation $\mu$ with parameter $k$, (recall the particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (4) and $S_{G}$ (2.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell^{2}}=\alpha_{1}\\ \dfrac{(r+y)k}{2\pi^{2}\ell^{2}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r+y}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell^{2}}=\alpha_{5}\\ \dfrac{ek^{2}}{8\pi^{2}\ell^{4}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell^{2}}\alpha_{3}=-2(1+\dfrac{y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{2ek}{(r+y)\ell^{2}}\alpha_{3}\\ r+y=\dfrac{2\pi^{2}\ell^{2}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (6)
@@
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{k(r+2y)}{4\pi^{2}\ell^{2}}=\dfrac{2k(r+2y)}{(r+y)\ell^{2}}\alpha_{3}=-(1+\dfrac{2y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r+2y}\alpha_{2}\end{cases}$ (7)
and the resulting action reads:
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(\bar{\Omega}_{\mathfrak{h}})$ (8)
$=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell^{2}}\big{(}\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(r+2y)R^{ab}\wedge\beta_{a}\wedge\beta_{b}\big{)}}^{Holst}+\overbrace{\dfrac{ek^{2}}{8\pi^{2}\ell^{4}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\quad+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}.$ (9)
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}(R^{ab}+\dfrac{2k}{\ell^{2}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{(r+y)k}{\pi^{2}\ell^{2}}R^{ab}\wedge\beta_{b}\delta_{ac}=\tau_{c}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta\beta^{c}}$ (10)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{(\tfrac{r}{2}+y)k}{\pi^{2}\ell^{2}}T_{a}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta A^{ab}}$ (11)
@@
For $r=y=0$, setting $e=\dfrac{4\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$ and applying the Hodge star * to the first equation (10) gives us the Einstein equations with bare cosmological constant $\Lambda_{0}$ while the other (11) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity:
@@
where $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell^{2}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell^{2}}-\dfrac{2\pi^{2}\ell^{2}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ is the effective cosmological constant, and $\Lambda_{\text{vac}}$ is the vacuum contribution.
@@
  • If $\Lambda_{\text{vac}}>\Lambda_{exp}$, then, agreement with experiment could be attained by fixing the appropriate value of $\Lambda_{0}=-\dfrac{6k}{\ell^{2}}$ with $k>0$, in particular, for $k=1$ $\Leftrightarrow G=SO(3,2)$ (AdS spacetime), thus potentially allowing the use of AdS/CFT correspondence [aharony__2008].
@@
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this either implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkoswki spacetime or $\Lambda_{0}=-\dfrac{6k}{\ell^{2}}\to 0$. However, for $k=0$ we do not retrieve Einstein’s equations since the action consists solely of the total derivatives corresponding to both the Euler and Pontrjagin densities of the curvature $R$.
@@
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $r=y=0$. However, even if they do not affect the classical equations of motion, the Euler and Pontrjagin densities of $R$ (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [8, p2].
@@
Using the same relation as in [8] we can identify the Barbero-Immirzi parameter $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{2e}{3r+4y}$, with our parameters. By replacing $e$, $y$ and $\ell$ in the action we obtain:
@@
$S_{G}[\varpi]$ $=\dfrac{3\pi}{2\Lambda_{0}G}\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+\dfrac{3}{4}(\dfrac{\pi}{\Lambda_{0}G\gamma}-r)\mathdutchcal{P}(\bar{\Omega}_{\mathfrak{h}})$ (16)
@@
$=\int_{\mathscr{M}}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{2}{\gamma}-\dfrac{2\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
@@
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{\underbrace{-(\dfrac{Gr}{\pi}+\dfrac{3}{2\Lambda_{0}\gamma})}_{\tfrac{2\vartheta}{\alpha_{1}}}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{4\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}.$

### v5

@@ 2.3 Associated action
The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the $\mathfrak{g}$-valued curvature $\bar{\Omega}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+\dfrac{2k}{\ell^{2}}\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}=K_{\mathfrak{g}}(\varphi,\varphi^{\prime})/3$, see (2.2).
By combining linearly the invariant polynomials $\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})$ and $\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}^{2})$ (corresponding respectively to deformations of the Euler and Pontrjagin numbers of $R$ with deformation $k\xi$) with the Pontrjagin number $\mathdutchcal{P}$ (see (7)) associated to $\bar{\Omega}$, one can build the action (for $e,r,y$ real numbers [^10] Complex numbers may be taken for example to describe a complex formulation [16, p.13] of Loop Quantum Gravity (LQG) with a complex Barbero-Immirzi parameter $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{e}{r+2y}$.):
$S_{G}[\varpi]$ $=r\mathdutchcal{P}(\overline{\Omega})+\int_{\mathscr{M}}(e\operatorname{Pf}(\dfrac{\overline{\Omega}_{\mathfrak{h}}}{2\pi})+y\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi}))$
$=r\mathdutchcal{P}(\bar{\Omega})+\int_{\mathscr{M}}\bigl{(}e\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})-\dfrac{y}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}\bar{\Omega}_{\mathfrak{h}})\bigr{)}=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}\left(\tfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi}\right)+rP(\bar{\Omega})-\dfrac{y}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
@@
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{4k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;b}d^{4}x+\dfrac{24k^{2}}{\ell^{4}}\sqrt{|\tilde{g}|}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)(R^{ab}\wedge R_{ab}+\dfrac{k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;d}\varepsilon^{cd}_{~{}~{}ab}d^{4}x)-\dfrac{2rk}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}$ (1)
@@
Using the isomorphism $\Phi_{\varepsilon}:\mathfrak{so}(n-1,1)\rightarrow\mathfrak{spin}(n-1,1)$ between the Lorentz Lie algebra $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$, [17, p.192], whose action on the Lie algebra generators of $\mathfrak{so}(n-1,1)$ is given by:
$\Phi_{\varepsilon}(J_{rs})=\dfrac{1}{4}\bar{e}_{a}J_{rs}e_{b}\gamma^{a}\gamma^{b}=\dfrac{1}{4}(J_{rs})_{ab}\gamma^{a}\gamma^{b}=\dfrac{1}{2}\eta_{ar}\eta_{bs}\gamma^{a}\gamma^{b}\ ,$ (2)
the covariant derivative associated to $\varpi_{\mathfrak{h}}$ on the space of spinors reads $D_{\mathfrak{h}}=d+\Phi_{\varepsilon}(\varpi_{\mathfrak{h}})=(\partial_{\mu}+\dfrac{1}{4}\varpi_{\mathfrak{h}\>\,ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$. One can then construct the following matter action: [^11] Since the metrics at the level of the Lie algebra and on $\mathscr{M}$ are always defined up to the factors $\kappa$ and $\zeta$, we set $\kappa^{2}\zeta=\tfrac{\ell^{4}}{216k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [17, p.197].
$\hskip-8.53581ptS_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=\int_{\mathscr{M}}\dfrac{36k^{2}\kappa\zeta^{2}}{\ell^{4}}h^{\varepsilon}\bigl{(}i\psi,\eta_{ab}\gamma^{a}\beta^{b}\wedge\tilde{*}D_{\mathfrak{h}}\psi\bigl{)}=\mathrm{Re}S_{D}$ (3)
where $\mathrm{Re}S_{D}$ corresponds to the real part of the Dirac action as described in [17, p.197] and $\tilde{*}$ is the Hodge star operator defined from $\tilde{g}$.
We are now in position to identify the different terms of $S_{G}[A,\beta]$ constructed above from the invariant polynomials and the Pontryagin characteristic class with the action presented in [16] [^12] This reference is preferably chosen due to the freedom of the parameters whereas the action given in [18][eqs (23-25)] is already more constrained. which, in our notation, reads:
$S^{\prime}[\varpi]$ $=\int_{\mathscr{M}}\Biggl{(}\biggl{(}\overbrace{\alpha_{1}\ell^{2}\underbrace{R^{ab}\wedge\star\xi_{ab}}_{Palatini}+\alpha_{2}\ell^{2}R^{ab}\wedge\xi_{ab}}^{Holst}+\overbrace{\alpha_{3}R^{ab}\wedge R_{ab}}^{Pontrjagin}$ (4)
$+\overbrace{\alpha_{4}R^{ab}\wedge\star R_{ab}}^{Euler}+\overbrace{\alpha_{5}\bigl{(}T^{a}\wedge T_{a}-\ell^{2}R^{ab}\wedge\xi_{ab}\bigr{)}}^{Nieh-Yan}+\overbrace{\alpha_{6}\ell^{4}\varepsilon_{abcd}\xi^{ab}\wedge\xi^{cd}}^{Cosmological~{}constant}\biggr{)}\Biggr{)}$
@@
In this case, from the first Bianchi identity $D\bar{\Omega}_{\mathfrak{h}}=0$, and by neglecting boundary terms, the equations of motion for $A$ and $\beta$ require respectively a null spin density and a null energy-momentum tensor as source terms and thus, turn out to be physically irrelevant.
While for general mutation $\mu$ with parameters $k,\ell$, (recall the particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (4) and $S_{G}$ (2.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{8\pi^{2}\ell^{2}}=\alpha_{1}\\ \dfrac{(r+y)k}{4\pi^{2}\ell^{2}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r+y}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell^{2}}=\alpha_{5}\\ \dfrac{ek^{2}}{32\pi^{2}\ell^{4}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{2k}{\ell^{2}}\alpha_{4}=\dfrac{4\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{2k}{\ell^{2}}\alpha_{3}=-(1+\dfrac{y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{ek}{(r+y)\ell^{2}}\alpha_{3}\\ r+y=\dfrac{4\pi^{2}\ell^{2}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (6)
@@
$\begin{cases}\alpha_{1}=\dfrac{2k}{\ell^{2}}\alpha_{4}=\dfrac{4\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{ky}{4\pi^{2}\ell^{2}}=\dfrac{2ky}{(r+y)\ell^{2}}\alpha_{3}=-\dfrac{y}{r}\alpha_{5}\\ \alpha_{1}=\dfrac{e}{2y}\alpha_{2}\end{cases}$ (7)
and the resulting deformed topological gauge action reads:
$S_{G}[\varpi]$ $=r\mathdutchcal{P}(\bar{\Omega})+\int_{\mathscr{M}}\bigl{(}e\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})-\dfrac{y}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\bar{\Omega}_{\mathfrak{h}})\bigr{)}$ (8)
$=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell^{2}}\big{(}\dfrac{e}{4}\underbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}_{Palatini}+yR^{ab}\wedge\beta_{a}\wedge\beta_{b}\big{)}}^{Holst}+\overbrace{\dfrac{ek^{2}}{32\pi^{2}\ell^{4}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Bare~{}Cosmological~{}constant}$
$\qquad\quad+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}.$
As already noticed in [16], there are two contributions to the Barbero-Immirzi parameter $\gamma$, which is expressed here as $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{e}{r+2y}$. One in front of $R^{ab}\wedge\beta_{a}\wedge\beta_{b}$ in the Holst term and the other one in the Nieh-Yan topological term.
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta\beta^{c}}$ $=\dfrac{ek}{8\pi^{2}\ell^{2}}(R^{ab}+\dfrac{k}{\ell^{2}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{yk}{2\pi^{2}\ell^{2}}R^{ab}\wedge\beta_{b}\eta_{ac}=\tau_{c}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta\beta^{c}}$ (9)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta A^{ab}}$ $=\dfrac{ek}{8\pi^{2}\ell^{2}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{yk}{2\pi^{2}\ell^{2}}T_{a}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta A^{ab}}$ (10)
@@
For $y=0$ and $\alpha_{1}=-\dfrac{1}{16\pi G}$, one identifies [^13] At this stage a remark is in order. Indeed, the $[\mathfrak{m},\mathfrak{m}]$ bracket occurring in the $\mathfrak{h}$-part of the curvature, see (12) which is related to the mutation parameters $k,\ell$ turns out to be related to the bare cosmological constant $\Lambda_{0}$. This gives a physical interpretation of the mutation. See Section 4 where the topological features of the theory will be linked to the vanishing of the bare cosmological constant $\Lambda_{0}$. $\ell^{2}=-\dfrac{3k}{\Lambda_{0}}$ and $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ such that applying the Hodge star * to the first equation (9) gives the Einstein equations with bare cosmological constant $\Lambda_{0}$ while (10) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity:
@@
One may notice that the previous identification allows to recover the value (up to the mutation parameter $k$ and sign convention) $e/(32\pi^{2})=-\ell^{2}/(64\pi Gk)$ of the parameter $\alpha$ of [19, Sec II] in front of the Euler density built out of $R$, leading to finite Noether charges and a regularized Euclidean action as mentioned in [20, 19].
@@
where $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{3k}{\ell^{2}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{3k}{\ell^{2}}-\dfrac{4\pi^{2}\ell^{2}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ is the effective cosmological constant, and $\Lambda_{\text{vac}}$ is the vacuum contribution.
@@
  • If $\Lambda_{\text{vac}}>\Lambda_{exp}$, then, agreement with experiment could be attained by fixing the appropriate value of $\Lambda_{0}=-\dfrac{3k}{\ell^{2}}$ with $k>0$, in particular, for $k=1$ $\Leftrightarrow G=SO(3,2)$ (AdS spacetime), thus potentially allowing the use of AdS/CFT correspondence [22].
@@
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this either implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkowski spacetime or $\Lambda_{0}=-\dfrac{3k}{\ell^{2}}\to 0$. In this limit, the contribution (12) to the $\mathfrak{h}$-part of the Cartan curvature becomes small and the theory can be considered as topological. This point will be discussed in detail in Section 4.
    However, for $k=0$ we do not retrieve Einstein’s equations since the action consists solely of the total derivatives corresponding to both the Euler and Pontrjagin densities of the curvature $R$.
@@
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $y=0$. Even if they do not affect the classical equations of motion, the Euler and Pontrjagin densities of $R$ as well as the Nieh-Yan term (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [16, p.2]. The Nieh-Yan term also gives a non trivial contribution to the Barbero-Immirzi parameter $\gamma=\dfrac{e}{r+2y}$ via the parameter $r$.
@@
Using the same relation as in [16] we can identify the Barbero-Immirzi parameter $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}=\dfrac{e}{r+2y}$, with our parameters. By replacing $e$, $y$ and $\ell$ in the action we obtain:
@@
$S_{G}[\varpi]$ $=r\mathdutchcal{P}(\overline{\Omega})+\int_{\mathscr{M}}(\dfrac{3\pi}{2\Lambda_{0}G}\operatorname{Pf}(\dfrac{\overline{\Omega}_{\mathfrak{h}}}{2\pi})+(\dfrac{3\pi}{4\Lambda_{0}G\gamma}-\dfrac{r}{2})\det(\mathds{1}+\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi}))$ (15)
@@
$=\int_{\mathscr{M}}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{2}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
@@
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-(\dfrac{2Gr}{\pi}+\dfrac{3}{\Lambda_{0}\gamma})R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{8\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}.$